# Training in reverse: How iteration order influences convergence and stability in deep learning

### v1

@@ A.1 Regression on synthetic datasets
In this experiment, to verify the increased stability of backward SGD over the standard forward version, we trained with both algorithms a neural network with 3 layers of 300 neurons with no regularization each to regress 100 points uniformly sampled from function graphs $D=\{(x_{i},y_{i}):\>y_{i}=f(x_{i})\;\textrm{with}\;x_{i}=-1+\frac{2i}{100},\,i=0,\dots,100\}$ using the functions $f(x)=x^{2}$ (Figure 4), $f(x)=\cos(10x)$ (Figure 5), and $f(x)=x^{3}$ (Figure 6). In all cases, we can observe a much higher stability for backward SGD over the forward version.
@@
Figure 5: Decreased variance and increased stability in train (left) and test (rigth) losses for backward SGD compared to forward SGD for all 5 seeds. The data was sampled from $f(x)=\cos(10x)$ and the training performed with batch size 1 and learning rate 0.02 for 1400 steps.
@@ A.2 MLP trained on Fashion-MNIST
In this experiment, to verify the increased stability of backward SGD over the standard forward version, we trained a MLP with 5 layers of 500 neurons each with no regularization using both forward and backward stochastic gradient descent with no regularization on the Fashion-MNIST dataset [45]. We repeated the experiment for 5 different random seeds. For all seeds, we used a learning rate of 0.001 and a batch size of 8. In Figure 7, we recorded the training loss at each gradient update for both the forward and backward iteration and plotted the 5 seeds separately (one seed for each row). For each seed, we observe that the training loss for the backward iterations is more stable and converges faster with less variability when compared to the forward iterations. We also observe stability improvements for all other learning curves for the backward trajectories.
@@ A.3 VGG-19 trained on CIFAR-10
In this experiment, to verify the increased stability of backward SGD over the standard forward version, we trained a VGG-19 model [46] using both forward and backward stochastic gradient descent with no regularization on the CIFAR-10 dataset [44]. We repeated the experiment for 5 different random seeds. In all seeds, we used a learning rate of 0.001 and a batch-size of 8. In Figure 8, we plot the training and test loss at each gradient update for both the forward and backward iteration. We plotted the 5 seeds separately, represented by each row. Note that for each seed, we observe that the training loss for the backward iterations is again more stable, converges faster, and has less variability than the forward iterations.
@@ A.4 ResNet-50 trained on CIFAR-100
In this experiment, to verify the increased stability of backward SGD over the standard forward version, we trained a ResNet-50 model [47] using both forward and backward stochastic gradient descent with no regularization on the CIFAR-100 dataset [44]. We repeated the experiment for 5 different random seeds. In all seeds, we used a learning rate of 0.001 and a batch-size of 16. In Figure 9, we plot the training and test loss at each gradient update for both the forward and backward iteration. We plotted the 5 seeds separately, represented by each row. Note that for each seed, we observe that the training loss for the backward iterations is again more stable, converges faster, and has less variability than the forward iterations.
@@
A.5 Backward stabilization after forward iterations
In this experiment, to verify the increased stability and convergence after turning on backward iterations after a number of forward iterations, we trained a MLP with 5 layers of 500 neurons each with stochastic gradient descent with no regularization to learn the 10 classes of Fashion-MNIST dataset [45]. We repeated the experiments for 5 seeds. We used a learning rate of 0.001 and a batch-size of 8. In Figure 10, we recorded the learning curves at each gradient update for both the forward iteration and the backward iteration switched on after step 1000. Each seed is plotted independently. We observe that the training loss for the backward iterations has again more stable convergence once backward iterations are switched on after step 1000 than the forward iterations.
@@ A.5 Backward stabilization after forward iterations
A.6 Continued stability of backward SGD throughout training.
In this experiment, to verify that the behavior of increased stability and convergence for backward iterates continues throughout training, we trained a ResNet-18 model on the CIFAR-10 dataset using the same hyperparameters as in Figure 2 (i.e., no regularization, a learning rate of 0.025 and batch size of 8) but training for 25000 steps instead of 2000. In Figure 11, we record the learning curves for both forward and backward SGD, performing model evaluation every 100 steps. Throughout training we continue to notice that the training loss (as well as all the other learning curves) for the backward iterations is again more stable, converges faster, and has less variability than the forward iterations. Note, here we only performed this experiment for one seed because because of the high computational requirements for backward SGD with so many training steps.
@@ Appendix B Additional seeds to plots in the main paper
B.1 All 5-seeds for ResNet-18 trained on CIFAR-10; cf. Figure 2
In Figure 12 below, we plot the learning curves for all five seeds as in Figure 2 of the main paper.
Figure 12: All 5-seeds plot of Figure 2.
B.2 All 5-seeds for MLP trained on FashionMNIST; cf. Figure 3
In Figure 13 below, we plot the learning curves for all five seeds as in Figure 3 of the main paper.
Figure 13: All 5-seeds plot of Figure 3.
@@ Appendix C Toward an approximate backward SGD
Backward SGD has increased stability and convergence over forward SGD but naive implementations are computationally intensive. In this section, we propose an approximation of the backward iterates by modifying the forward iterates with a Lie bracket term of order $\mathcal{O}(h^{2})$ in the learning rate (Theorem C.2). This comes at an extra cost, but which can be made smaller than evaluating the backward iterates from the initial point at every step (Corollary C.4). Our approximation is valid
@@
in Taylor’s series in the learning rate. This is what the next lemma gives. (We demonstrate the usefulness of such expansions Appendix D where we use Lemma C.1 to produce a beneficial second order regularizer that emulates an iteration order average.)
Observe also that Lemma C.1 tells us that all possible iteration orders coincide at order $h$ but start differing at order $\mathcal{O}(h^{2})$:
@@
We proceed by induction. For the base case, $k=1$, this is trivial. Suppose now that this is true for any composition of $k-1$ operators. By definition of $T_{i_{1}}$ we have that
@@
$X=\theta+h\sum_{l=2}^{k}V_{i_{l}}(\theta)+h^{2}\sum_{2\leq u<v\leq k}V_{i_{u}}^{\prime}(\theta)V_{i_{v}}(\theta)+\mathcal{O}(h^{3})$
Therefore, taking a Taylor series for the second term of (6), we obtain
$hV_{i_{1}}(X)=hV_{i_{1}}(\theta)+h^{2}\sum_{l=2}^{k}V_{i_{1}}^{\prime}(\theta)V_{i_{l}}(\theta)+\mathcal{O}(h^{3}).$
Summing up in (6) the expressions we have found for $X$ and $hV_{i_{1}}(X)$ above, we obtain that the composition $T_{i_{1}}\cdots T_{i_{k}}(\theta)$ has the form
@@
where the $[V_{i},V_{j}](\theta)=V_{i}^{\prime}(\theta)V_{j}(\theta)-V_{j}^{\prime}(\theta)V_{i}(\theta)$ is the Lie bracket between the vector fields $V_{i}$ and $V_{j}$.
@@
By Lemma C.1, we have that the backward iterate is
$T_{1}\cdots T_{n}=\operatorname{id}+h\sum_{l=1}^{n}V_{l}+h^{2}\sum_{1\leq u<v\leq n}V_{u}^{\prime}V_{v}+\mathcal{O}(h^{3})$
@@
Consider the SGD operators $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$ obtained by taking the gradient of a loss function on a batch $B_{i}$ of data at step $i$. We denote by $\theta_{n}=T_{n}\cdots T_{1}(\theta_{0})=\theta_{n-1}-h\nabla L_{n}(\theta_{n-1})$ the forward SGD iterate starting at initial point $\theta_{0}$ and by $\theta_{n}^{B}=T_{1}\cdots T_{n}(\theta_{0})$ the corresponding backward iterate starting at the same initial point. Motivated by Theorem C.2, we introduce the approximate backward iterate $\tilde{\theta}_{n}$ as follows:
@@
Theorem C.2 tells us that the true backward iterate and its approximation are within $\mathcal{O}(h^{3})$ of each other:
@@
By Definition C.3 of the approximate backward iterates we have that $\tilde{\theta}_{n}=\theta_{n}+h^{2}C_{n}$ with
@@ Appendix D Approximate implicit regularization of smaller batches
As shown in [17, 48, 15, 16] for instance, small batches have an implicit regularization effect, producing solutions with higher test accuracy as the batch size decreases. In this section, we show that approximations of the type given by Lemma C.1 are useful to understand this implicit regularization effect. We also see how we can go beyond this small batch effect and produce explicit regularizers with even higher test performance based on the idea of iteration order average on the small batches, leveraging Lemma C.1.
Throughout the section we will write $V_{i}(\theta)$ for the negative batch loss gradient $-\nabla L_{i}(\theta)$ computed on batch $B_{i}$ in order to keep the notation simple.
@@ D.1 The effect of small batches
The idea is to use Lemma C.1 to understand the implicit regularization effect of small batches, very much in line with the findings in [15]. First, consider two settings. In the first setting, we perform a single gradient update
$T_{\textrm{large}}(\theta)=\theta+hV_{B}(\theta).$
@@
Then Lemma C.1 gives us immediately a second order approximation for the second setting:
@@
So we can attribute the second order regularization effect of small batch training to this additional term, in line with the computations in [15].
@@ D.2 Explicit regularization through iteration order averages
where $\mathrm{Perm}_{c}$ is the permutations of $c$ elements. Again Lemma C.1 gives us the second order approximation for this new update rule:
@@ D.2.1 Experiments
We trained a MLP with 5 layers of 500 neurons each with stochastic gradient descent with no regularization to learn the 10 classes of Fashion-MNIST dataset [45]. We used a learning rate of $h=0.001$ (and $h^{\prime}=h/c$), and a batch-size of $|B|=2048$§§Or the closest integer divisible by $c$; e.g., $2049$ for $c=3$.. We examine three different shrinking factors $c\in\{2,3,4\}$. In each setting, we test the values of $\lambda\in\{0.5h^{2},h^{2},2h^{2},4h^{2}\}$. The following figures present the train and test loss over training compared to training with large-batch $|B|=2048$ and learning rate $h^{\prime}=h/c$, and training with small-batch $2048/c$ with learning rate $h^{\prime}$. The experiment shows that our method can outperform vanilla training both for large and small batches with minimal tuning of $\lambda$ for this specific experiment.
@@
Appendix E Proof of Theorem 2.6
@@ E.1 Definitions
This can be thought of as a sort of point-wise convergence. A weaker notion of convergence which is implied by almost sure convergence (see Theorem 2.7 (i) in [49]) is convergence in probability. The sequence $X_{n}$ of random variables converges to the random variable $X_{\infty}$ in probability if for every $\varepsilon>0$ the following holds
@@
That is, for large $n$ it becomes more and more likely that $X_{n}$ is close to $X_{\infty}$. The weakest notion of convergence is convergence in distribution, which is implied by convergence in probability (see Theorem 2.7 (ii) in [49]). There are many equivalent ways to define this, but the most concrete way is the following: Let $F_{n}(x)=\mathbb{P}(X_{n}\leq x)$ be the cumulative distribution function for $X_{n}$ and let $F_{\infty}(x)$ be the corresponding cumulative distribution function for $X_{\infty}$, then we say that the sequence $X_{n}$ converges in distribution to $X_{\infty}$ if for every continuity point $x$ of the cumulative distribution function $F_{\infty}$ we have the following

### v2

@@ Appendix A Additional Experiments
Note on plotting multiple seeds: We are interested in the variability per realization of the backward and forward trajectories. The backward trajectories are more stable individually along their own paths, but these paths can be very different from seed to seed (because of the convergence toward different points). Therefore, the phenomenon is much clearer when the seeds are plotted individually rather than averaged, which creates artificially more variability in the backward trajectories as there really is on each individual realization. This is why we are reporting the various seed plots individually and not as a single averaged plot with error bars. Note that the increased stability and point convergence is visible in each of the seeds.
@@ A.1 Regression on synthetic datasets
In this experiment, to verify the increased stability of backward SGD over the standard forward version, we trained with both algorithms a neural network with 3 layers of 300 neurons with no regularization each to regress 100 points uniformly sampled from function graphs $D=\{(x_{i},y_{i}):\>y_{i}=f(x_{i})\;\textrm{with}\;x_{i}=-1+\frac{2i}{100},\,i=0,\dots,100\}$ using the functions $f(x)=x^{2}$ (Figure˜4), $f(x)=\cos(10x)$ (Figure˜5), and $f(x)=x^{3}$ (Figure˜6). In all cases, we can observe a much higher stability for backward SGD over the forward version.
@@
Figure 5: Decreased variance and increased stability in train (left) and test (right) losses for backward SGD compared to forward SGD for all 5 seeds. The data was sampled from $f(x)=\cos(10x)$ and the training performed with batch size 1 and learning rate 0.02 for 1400 steps.
@@ A.2 MLP trained on Fashion-MNIST
In this experiment, to verify the increased stability of backward SGD over the standard forward version, we trained a MLP with 5 layers of 500 neurons each with no regularization using both forward and backward stochastic gradient descent with no regularization on the Fashion-MNIST dataset Xiao et al. (2017). We repeated the experiment for 5 different random seeds. For all seeds, we used a learning rate of 0.001 and a batch size of 8. In Figure˜7, we recorded the training loss at each gradient update for both the forward end backward iteration and plotted the 5 seeds separately (one seed for each row). For each seed, we observe that the training loss for the backward iterations is more stable and converges faster with less variability when compared to the forward iterations. We also observe stability improvements for all other learning curves for the backward trajectories.
@@ A.3 VGG-19 trained on CIFAR-10
In this experiment, to verify the increased stability of backward SGD over the standard forward version, we trained a VGG-19 model Simonyan and Zisserman (2014) using both forward and backward stochastic gradient descent with no regularization on the CIFAR-10 dataset Krizhevsky (2009). We repeated the experiment for 5 different random seeds. In all seeds, we used a learning rate of 0.001 and a batch-size of 8. In Figure˜8, we plot the training and test loss at each gradient update for both the forward and backward iteration. We plotted the 5 seeds separately, represented by each row. Note that for each seed, we observe that the training loss for the backward iterations is again more stable, converges faster, and has less variability than the forward iterations.
@@ A.4 ResNet-50 trained on CIFAR-100
In this experiment, to verify the increased stability of backward SGD over the standard forward version, we trained a ResNet-50 model He et al. (2016) using both forward and backward stochastic gradient descent with no regularization on the CIFAR-100 dataset Krizhevsky (2009). We repeated the experiment for 5 different random seeds. In all seeds, we used a learning rate of 0.001 and a batch-size of 16. In Figure˜9, we plot the training and test loss at each gradient update for both the forward and backward iteration. We plotted the 5 seeds separately, represented by each row. Note that for each seed, we observe that the training loss for the backward iterations is again more stable, converges faster, and has less variability than the forward iterations.
@@
A.5 ResNet-18 trained on CIFAR-10 with AdamW
In this experiment we verify improved stability of backward iterates for a base optimizer other than vanilla SGD. AdamW Loshchilov and Hutter (2017) combines momentum from Adam with weight decay regularization and improves upon the standard Adam by preventing the adaptive learning rates from distorting the intended regularization strength.
Here, we use AdamW for both forward and backward gradient updates, only altering the order in which batches are consumed. We train a ResNet-18 model He et al. (2016) and do not apply any regularization on the CIFAR-10 dataset Krizhevsky (2009). We repeated the experiment for 5 different random seeds. In all seeds, we used a learning rate of 0.00025 and a batch-size of 8. In Figure˜10, we plot the training and test loss at each gradient update for both the forward and backward iteration. We plotted the 5 seeds separately, represented by each row. Note that for each seed, we observe that the training loss for the backward iterations is again more stable, converges faster, and has less variability than the forward iterations.
Figure 10: ResNet-18 trained on CIFAR-10 with AdamW. Training with backward AdamW significantly reduces the variance and increases the stability of both the train and test loss as well as the train and test and accuracy compared to forward AdamW. This behavior is consistent across all seeds.
A.6 Backward stabilization after forward iterations
In this experiment, to verify the increased stability and convergence after turning on backward iterations after a number of forward iterations, we trained a MLP with 5 layers of 500 neurons each with stochastic gradient descent with no regularization to learn the 10 classes of Fashion-MNIST dataset Xiao et al. (2017). We repeated the experiments for 5 seeds. We used a learning rate of 0.001 and a batch-size of 8. In Figure˜11, we recorded the learning curves at each gradient update for both the forward iteration and the backward iteration switched on after step 1000. Each seed is plotted independently. We observe that the training loss for the backward iterations has again more stable convergence once backward iterations are switched on after step 1000 than the forward iterations.
@@ A.6 Backward stabilization after forward iterations
A.7 Continued stability of backward SGD throughout training.
In this experiment, to verify that the behavior of increased stability and convergence for backward iterates continues throughout training, we trained a ResNet-18 model on the CIFAR-10 dataset using the same hyperparameters as in Figure˜2 (i.e., no regularization, a learning rate of 0.025 and batch size of 8) but training for 25000 steps instead of 2000. In Figure˜12, we record the learning curves for both forward and backward SGD, performing model evaluation every 100 steps. Throughout training, we continue to notice that the training loss (as well as all the other learning curves) for the backward iterations is again more stable, converges faster, and has less variability than the forward iterations. Note, here we only performed this experiment for one seed because because of the high computational requirements for backward SGD with so many training steps.
@@ Appendix B Additional seeds to plots in the main paper
B.1 All 5-seeds for ResNet-18 trained on CIFAR-10; cf. Figure˜2
In Figure˜13 below, we plot the learning curves for all five seeds as in Figure˜2 of the main paper.
Figure 13: All 5-seeds plot of Figure˜2.
B.2 All 5-seeds for MLP trained on FashionMNIST; cf. Figure˜3
In Figure˜14 below, we plot the learning curves for all five seeds as in Figure˜3 of the main paper.
Figure 14: All 5-seeds plot of Figure˜3
@@ Appendix C Toward an approximate backward SGD
Backward SGD has increased stability and convergence over forward SGD but naive implementations are computationally intensive. In this section, we propose an approximation of the backward iterates by modifying the forward iterates with a Lie bracket term of order $\mathcal{O}(h^{2})$ in the learning rate (Theorem˜C.2). This comes at an extra cost, but which can be made smaller than evaluating the backward iterates from the initial point at every step (Corollary˜C.4). Our approximation is valid
@@
in Taylor’s series in the learning rate. This is what the next lemma gives. (We demonstrate the usefulness of such expansions Appendix˜D where we use Lemma˜C.1 to produce a beneficial second order regularizer that emulates an iteration order average.)
Observe also that Lemma˜C.1 tells us that all possible iteration orders coincide at order $h$ but start differing at order $\mathcal{O}(h^{2})$:
@@
We proceed by induction. For the base case, $k=1$, this is trivial. Suppose now that this is true for any composition of k1k-1 operators. By definition of $T_{i_{1}}$ we have that
@@
$X=\theta+h\sum_{l=2}^{k}V_{i_{l}}(\theta)+h^{2}\sum_{2\leq u<v\leq k}V_{i_{u}}^{\prime}(\theta)V_{i_{v}}(\theta)+\mathcal{O}(h^{3})$ (8)
Therefore, taking a Taylor series for the second term of ˜7, we obtain
$hV_{i_{1}}(X)=hV_{i_{1}}(\theta)+h^{2}\sum_{l=2}^{k}V_{i_{1}}^{\prime}(\theta)V_{i_{l}}(\theta)+\mathcal{O}(h^{3}).$ (9)
Summing up in ˜7 the expressions we have found for $X$ and $hV_{i_{1}}(X)$ above, we obtain that the composition $T_{i_{1}}\cdots T_{i_{k}}(\theta)$ has the form
@@
where the $V_{i},V_{j}](\theta)=V_{i}^{\prime}(\theta)V_{j}(\theta)-V_{j}^{\prime}(\theta)V_{i}(\theta)$ is the Lie bracket between the vector fields $V_{i}$ and $V_{j}$.
@@
By Lemma˜C.1, we have that the backward iterate is
$T_{1}\cdots T_{n}=1+h\sum_{l=1}^{n}V_{l}+h^{2}\sum_{1\leq u<v\leq n}V_{u}^{\prime}V_{v}+\mathcal{O}(h^{3})$
@@
Consider the SGD operators $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$ obtained by taking the gradient of a loss function on a batch $B_{i}$ of data at step $i$. We denote by $\theta_{n}=T_{n}\cdots T_{1}(\theta_{0})=\theta_{n-1}-h\nabla L_{n}(\theta_{n-1})$ the forward SGD iterate starting at initial point $\theta_{0}$ and by $\theta_{n}^{B}=T_{1}\cdots T_{n}(\theta_{0})$ the corresponding backward iterate starting at the same initial point. Motivated by Theorem˜C.2, we introduce the approximate backward iterate $\tilde{\theta}_{n}$ as follows:
@@
Theorem˜C.2 tells us that the true backward iterate and its approximation are within $\mathcal{O}(h^{3})$ of each other:
@@
By Definition˜C.3 of the approximate backward iterates we have that $\tilde{\theta}_{n}=\theta_{n}+h^{2}C_{n}$ with
@@ Appendix D Approximate implicit regularization of smaller batches
As shown in Keskar et al. (2017); Smith et al. (2017; 2021); Dherin et al. (2022) for instance, small batches have an implicit regularization effect, producing solutions with higher test accuracy as the batch size decreases. In this section, we show that approximations of the type given by Lemma˜C.1 are useful to understand this implicit regularization effect. We also see how we can go beyond this small batch effect and produce explicit regularizers with even higher test performance based on the idea of iteration order average on the small batches, leveraging Lemma˜C.1.
Throughout the section we will write $V_{i}(\theta)$ for the negative batch loss gradient $\nabla L_{i}(\theta)$ computed on batch $B_{i}$ in order to keep the notation simple.
@@ D.1 The effect of small batches
The idea is to use Lemma˜C.1 to understand the implicit regularization effect of small batches, very much in line with the findings in Smith et al. (2021). First, consider two settings. In the first setting, we perform a single gradient update
$T_{\textrm{large}}(\theta)=\theta+hV_{B}(\theta),$
@@
Then Lemma˜C.1 gives us immediately a second order approximation for the second setting:
@@
So we can attribute the second order regularization effect of small batch training to this additional term, in line with the computations in Smith et al. (2021).
@@ D.2 Explicit regularization through iteration order averages
where $\mathrm{Perm}_{c}$ is the permutations of $c$ elements. Again Lemma˜C.1 gives us the second order approximation for this new update rule:
@@ D.2.1 Experiments
We trained a MLP with 5 layers of 500 neurons each with stochastic gradient descent with no regularization to learn the 10 classes of Fashion-MNIST dataset Xiao et al. (2017). We used a learning rate of $h=0.001$ (and $h^{\prime}=h/c$), and a batch-size of |B|=2048|B|=2048222Or the closest integer divisible by $c$; e.g., $2049$ for $c=3$.. We examine three different shrinking factors $c\in\{2,3,4\}$. In each setting, we test the values of $\lambda\in\{0.5h^{2},h^{2},2h^{2},4h^{2}\}$. The following figures present the train and test loss over training compared to training with large-batch $|B|=2048$ and learning rate $h^{\prime}=h/c$, and training with small-batch $2048/c$ with learning rate $h^{\prime}$. The experiment shows that our method can outperform vanilla training both for large and small batches with minimal tuning of $\lambda$ for this specific experiment.
@@
Appendix E Proof of Theorem˜2.6
@@ E.1 Definitions
This can be thought of as a sort of point-wise convergence. A weaker notion of convergence which is implied by almost sure convergence (see Theorem 2.7 (i) in Van der Vaart (2000)) is convergence in probability. The sequence $X_{n}$ of random variables converges to the random variable $X_{\infty}$ in probability if for every $\varepsilon>0$ the following holds
@@
That is, for large $n$ it becomes more and more likely that $X_{n}$ is close to $X_{\infty}$. The weakest notion of convergence is convergence in distribution, which is implied by convergence in probability (see Theorem 2.7 (ii) in Van der Vaart (2000)). There are many equivalent ways to define this, but the most concrete way is the following: Let $F_{n}(x)=\mathbb{P}(X_{n}\leq x)$ be the cumulative distribution function for $X_{n}$ and let $F_{\infty}(x)$ be the corresponding cumulative distribution function for $X_{\infty}$, then we say that the sequence $X_{n}$ converges in distribution to $X_{\infty}$ if for every continuity point $x$ of the cumulative distribution function $F_{\infty}$ we have the following